# Unisolver: PDE-Conditional Transformers Are Universal PDE Solvers

### v3

@@ F.1 Metrics
We can calculate the relative L2 error of model prediction $\hat{u}$ and ground truth $u$ as follows:
$\frac{\|u-\hat{u}\|_{L^{2}}}{\|u\|_{L^{2}}}.$ (17)
where $\|u-\hat{u}\|_{L^{2}}$ is the $L^{2}$-distance between the predicted solution $\hat{u}$ and the ground-truth solution $u$, and $\|u\|_{L^{2}}$ is the $L^{2}$-norm of the ground-truth solution.
@@
F.2 Implementations for Each Benchmark
@@
As outlined in Section 4, all the baseline models are trained under the same training strategy. Specifically, all the models are trained for 300 epochs using the relative L2 loss and the ADAM optimizer (Kingma & Ba, 2015) with an initial learning rate of $0.0005$, along with a cosine annealing learning rate scheduler (Loshchilov & Hutter, 2016). The batch size is set to 60. After the training process, we use the checkpoint from the last epoch to evaluate the model performance.
We also provide the detailed model architecture hyperparameters in Table 8. We configure each model to align their model parameter numbers to ensure a fair comparison. Note that for MPP (2023), We utilize the parameter configuration of the tiny version containing approximately five million trainable parameters, which is comparable to Unisolver and other baselines.
The varying PDE components in this benchmark include the viscosity coefficient and the external force. This physics information is provided to each baseline in an explicit or implicit way to ensure a fair comparison. For FNO (2021a), ViT (2020), Factformer (2023b), and MPP (2023), we explicitly concatenate the viscosity coefficient and the external force to the model input along the channel dimension to ensure a fair comparison. As the viscosity coefficient is essentially a scalar, we repeat it along the spatial dimensions and then perform the channel-concatenating process. ICON (2023) is a special baseline which takes prompting trajectories as additional inputs to implicitly extract the physics information. Consequently, instead of providing the PDE components, we augment the input to ICON with five additional prompting trajectories with the same viscosity and external force as the target trajectory. Note that ICON also needs additional prompting trajectories when conducting evaluation. For PINO (2021b), we follow the experiment setting in the original paper and train the model with physics-informed loss as a soft regularization. The proportion of physics-informed loss with regard to data loss is set to 0.1.
@@
We compare Unisolver with PDEformer-L in the 1D time-dependent PDEs benchmark, evaluating their in-distribution test and zero-shot generalization performance. We also report the model performance after fine-tuning in Appendix A. The pre-training and fine-tuing configurations for Unisolver and the fine-tuing configurations for PDEformer are listed in Table 9.
After the pre-training process, we use the checkpoint from the last epoch to evaluate the model performance for the in-distribution test and zero-shot generalization test in Section 4.2. For fine-tuning tasks, we utilize the fine-tuning script provided in the original repository of PDEformer and set the finetuning epochs to 100 for a fair comparison.
The model we use to compare with PDEformer-L with contains 19M trainable parameters, which is comparable to the 22M parameters of PDEformer-L. The model scalability experiments in Section 4.4 also show model configurations with different number of trainable parameters. We progressively increase the Unisolver parameter from 3M to 63M, thus resulting in 4 different model configurations. We present the detailed configurations of these models in Table 10. Note that in this benchmark, we utilize an adapted version PolyINR (Singh et al., 2023) to decode the encoder output from the Transformer backbone.
@@
We compare Unisolver with DPOT-S with comparable model parameters in the 2D mixed PDEs benchmark. The training hyperparameter and model configurations are presented in Table 11.
This benchmark includes multiple diverse PDEs, each including its unique PDE components as illustrated in Appendix E. For example, the viscosity coefficient is the varying PDE components in the FNO-$\nu$ benchmark, while the shallow-water equation does not include this PDE component. Therefore, we must notice Unisolver whether a PDE component exists in a certain benchmark. To do so, Specifically, we introduce a binary masking channel to represent the existence of a certain PDE component. For example, when a PDE component exists in a benchmark, we concatenate an “1” with this component, indicating that this component is a valid one. When this PDE component does not exist in a benchmark, we concatenate an “0” with it, indicating that this is an invalid one.
@@ F.3 Details of the LLM embeddings
Note that the pre-training dataset of PDEformer (Ye et al., 2024) contains the PDE family following the formulation:
@@
Appendix G Additional Analyses
G.1 Unseen Viscosity and Unseen External Force on HeterNS
In addition to Tables 3 and 4, we further assess Unisolver’s generalization on HeterNS compared to other baselines under more challenging conditions, where both the viscosity coefficient and external force are unseen. Specifically, we generate nine different component pairs $(\nu,\omega)$, each with 200 testing samples. Notably, one case features $\omega$ = 6 , which significantly exceeds the maximum value of $\omega=3$ used during training, making it particularly difficult. The full results are presented in Table 13. Unisolver consistently outperforms all baselines, especially in the most challenging case with $\omega=6$, with a relative promotion of 37.1%.
@@ G.1 Unseen Viscosity and Unseen External Force on HeterNS
G.2 More ablation studies
@@
As shown in Tables 14 and 15, removing the information leads to a significant drop in performance compared to vanilla Unisolver, and concatenating the information directly also results in a huge decline. It is worth noting that the absence of external force information or its improper use (e.g. via direct concatenation) significantly degrades performance even in zero-shot viscosity generalization tasks, and vice versa, further highlighting the importance of including complete PDE components.
@@
G.3 Equations Generalization via fine-tuning
We design a new equation generalization scenario based on the 1D time-dependent PDEs benchmark. As stated in Appendix E, the general equation formulations used in this benchmark include two polynomials, $f_{0}$ and $f_{1}$, both with a maximum order of 3. We pretrain Unisolver on 50,000 samples of PDEs with polynomial orders of up to 2, and then fine-tune it for 200 epochs on PDEs with polynomial orders of 3. The fine-tuned model is compared against the same model trained from scratch for 500 epochs, with relative L2 error reported in Table 16. Results indicate that Unisolver pretrained on equations of polynomial order up to 2 can be efficiently fine-tuned to handle equations of polynomial order 3. Unisolver demonstrates strong generalization capabilities to unseen PDEs, significantly reducing the need for large training datasets when addressing new equations.
         Finetuning Examples          5000          10000          20000
         Unisolver-from-scratch-500          0.3308          0.1913          0.1327
         Unisolver-fine-tune-200          0.1624          0.1036          0.0891
         Promotion          50.9%          45.8%          32.9%
G.4 Full Scalability
As a supplement to Figure 8 in the main text, we also conduct experiments on different zero-shot generalization tasks from (Takamoto et al., 2022) and record the concrete data in Table 17 for clarity.
@@ G.4 Full Scalability
G.5 Standard Deviations
We repeat the experiments three times on the HeterNS benchmark and provide standard deviations here. As shown in Table 18-19, Unisolver surpasses the previous state-of-the-art models with high confidence. Note that we compare Unisolver with the second-best model, which is a strong baseline as it is not achieved by a single model. The results demonstrate that Unisolver significantly outperforms baseline models, with the second-best result falling more than three standard deviations behind, except in the case of viscosity $\nu=8e-4$.
@@ G.5 Standard Deviations
G.6 Detailed Compute Resources
@@
Appendix H Limitations and Future Work
This paper presents Unisolver to solve PDEs under universal PDE components, which achieves impressive performance supported by extensive analyses and visualizations. However, our method is currently limited to grid data due to the patchifying process during the embedding of point-wise components. Actually, this limitation is shared in all the generalizable PDE solvers, such as MPP (McCabe et al., 2023), Poseidon (Herde et al., 2024), PDEformer (Ye et al., 2024) and DPOT (Hao et al., 2024). One fundamental reason is the lack of suitable and large-scale irregular-mesh PDE datasets, which will require extremely high computation costs for generation and massive resources for collection. Since our primary focus in this paper is on the study of model architecture design and generalization capabilities, we would like to leave the irregular-mesh PDE dataset as a future work. Also, the capability to handle irregular meshes of Unisolver can be achieved by replacing the canonical Transformer with the latest geometry-general PDE solver: Transolver (Wu et al., 2024).

### v4

@@ H.1 Loss and Metrics
We can calculate the relative L2 distance between ground truth $u$ and model prediction $\hat{u}$ as follows:
[…]
where $\|u-\hat{u}\|_{L^{2}}$ is the $L^{2}$-distance between the predicted solution $\hat{u}$ and the ground-truth solution $u$, and $\|u\|_{L^{2}}$ is the $L^{2}$-norm of the ground-truth solution. Relative L2 is used as both training loss and evaluation metric.
@@
H.2 Implementations for Each Benchmark
@@
As outlined in Section 4, all the baseline models are trained under the same training strategy. We train the model using one-step predictions and test the model in an autoregressive manner. Specifically, all the models are trained for 300 epochs using the relative L2 loss and the ADAM optimizer [17] with an initial learning rate of $0.0005$, along with a cosine annealing learning rate scheduler [26]. The batch size is set to 60. After the training process, we use the checkpoint from the last epoch to evaluate the model performance.
We also provide the detailed model architecture hyperparameters in Table 12. We configure each model to align their model parameter numbers to ensure a fair comparison. Note that for MPP [2023], We utilize the parameter configuration of the tiny version containing approximately five million trainable parameters, which is comparable to Unisolver and other baselines.
The varying PDE components in this benchmark include the viscosity coefficient and external force. This physics information is provided to each baseline in an explicit or implicit way to ensure a fair comparison. For FNO [2021a], ViT [2020], FactFormer [2023b], and MPP [2023], we explicitly concatenate the viscosity coefficient and the external force to the model input along the channel dimension to ensure a fair comparison. As the viscosity coefficient is essentially a scalar, we repeat it along the spatial dimensions and then perform the channel-concatenating process. ICON [2023] is a special baseline which takes prompting trajectories as additional inputs to implicitly extract the physics information. Consequently, instead of providing the PDE components, we augment the input to ICON with five additional prompting trajectories with the same viscosity and external force as the target trajectory. Note that ICON also needs additional prompting trajectories when conducting evaluation. For PINO [2021b], we follow the experiment setting in the original paper and train the model with physics-informed loss as a soft regularization. The proportion of physics-informed loss with regard to data loss is set to 0.1.
@@
We compare Unisolver with PDEformer-L in the 1D time-dependent PDEs benchmark, evaluating their in-distribution test and zero-shot generalization performance. We also report the model performance after fine-tuning in Appendix B. The pre-training and fine-tuing configurations for Unisolver and the fine-tuing configurations for PDEformer are listed in Table 13.
Following PDEformer’s training strategies, we train the model to predict the solution at specific spatial-temporal coordinates through an INR. After the pre-training process, we use the checkpoint from the last epoch to evaluate the model performance for the in-distribution test and zero-shot generalization test in Section 4.2. For fine-tuning tasks, we utilize the fine-tuning script provided in the original repository of PDEformer and set the finetuning epochs to 100 for a fair comparison.
The model we use to compare with PDEformer-L contains 19M trainable parameters, which is comparable to the 22M parameters of PDEformer-L. The model scalability experiments in Section 4.4 also show model configurations with different number of trainable parameters. We progressively increase the Unisolver parameter from 3M to 63M, resulting in 4 different model configurations. We present the detailed configurations of these models in Table 14. Note that in this benchmark, we utilize an adapted version PolyINR [35] to decode the encoder output from the Transformer backbone.
@@
We compare Unisolver with DPOT-S with comparable model parameters in the 2D mixed PDEs benchmark. The training hyperparameter and model configurations are presented in Table 15. Similar to the HeterNS benchmark, We train the model using one-step predictions and test the model in an autoregressive manner.
This benchmark includes multiple diverse PDEs, each including its unique PDE components as illustrated in Appendix G. For example, the viscosity coefficient is the varying PDE components in the FNO-$\nu$ benchmark, while the shallow-water equation does not include this PDE component. Therefore, we must notice Unisolver whether a PDE component exists in a certain benchmark. To do so, Specifically, we introduce a binary masking channel to represent the existence of a certain PDE component. For example, when a PDE component exists in a benchmark, we concatenate an “1” with this component, indicating that this component is a valid one. When this PDE component does not exist in a benchmark, we concatenate an “0” with it, indicating that this is an invalid one. While the LLM embedding can provide some indication of this information, it does not serve as the input to the encoders of other components. This binary mask, however, aids the encoders’ learning and further clarifies the information without introducing significant computational overhead.
@@ H.3 Details of the LLM embeddings
Note that the pre-training dataset of PDEformer [45] contains the PDE family following the formulation:
@@
Appendix I Additional Analyses
I.1 Unseen Viscosity and Unseen External Force on HeterNS
In addition to Tables 10 and  11, we further assess Unisolver’s generalization on HeterNS compared to other baselines under more challenging conditions, where both the viscosity coefficient and external force are unseen. Specifically, we generate nine different component pairs $(\nu,\omega)$, each with 200 testing samples. Notably, one case features $\omega$ = 6 , which significantly exceeds the maximum value of $\omega=3$ used during training, making it particularly difficult. The full results are presented in Table 17. Unisolver consistently outperforms all baselines, especially in the most challenging case with $\omega=6$, with a relative promotion of 37.1%.
@@ I.1 Unseen Viscosity and Unseen External Force on HeterNS
I.2 More Ablation Studies on PDE Components and Conditional Modeling
@@
As shown in Tables 18 and 19, removing the information leads to a significant drop in performance compared to vanilla Unisolver, and concatenating the information directly also results in a huge decline. It is worth noting that the absence of external force information or its improper use (e.g. via direct concatenation) significantly degrades performance even in zero-shot viscosity generalization tasks, and vice versa, further highlighting the importance of including complete PDE components.
@@
I.3 Long Trajectory Prediction
We extend the temporal evolution steps of HeterNS to 30 steps, corresponding to 30 seconds of complex fluid dynamics, and report the zero-shot performance comparison between Unisolver and the baselines in the Table 20. We present the performance on the subdataset with a viscosity coefficient of $\nu=1\times 10^{-5}$ and a force coefficient of $\omega=2$. This is a particularly challenging task, as these models have never seen such long trajectories in the training data (at most 20 seconds). Despite this, Unisolver still achieves the best performance compared with the top three baselines.
Table 20: Zero-shot performance comparison (relative L2) with top three baselines
         Unisolver          FNO          ViT          FactFormer
         Relative L2          0.1956          0.3105          0.2527          0.2962
I.4 Full Scalability
As a supplement to Figure 10 in the main text, we also conduct experiments on different zero-shot generalization tasks from [37] and record the concrete data in Table 21 for clarity.
@@ I.4 Full Scalability
I.5 Efficiency Analysis
We provide the inference time and memory consumption for each model to predict a single frame on the HeterNS, along with the calculation time and memory consumption of the numerical solver, which is a pseudo-spectral solver adopted by FNO [19] used to generate the HeterNS dataset, to calculate the next frame, as summarized in the Table 22. The results are measured on an A100 GPU with a batch size of 1. Unisolver demonstrates comparable inference speed to FNO, while consuming less memory. Besides, all neural PDE solvers are approximately 1,000 times faster than the numerical solver, highlighting their potential as efficient surrogate models.
Table 22: Efficiency Analysis. The inference (calculation) time and memory consumption for each model and numerical solver to predict a single frame on the HeterNS dataset.
FNO PINO ViT FactFormer ICON MPP Unisolver
Numerical Solver
Average Inference 0.0042 0.0042 0.0045 0.0103 0.0057 0.0120 0.0054 7.26
(Calculation) Time / s
Average Memory 730 730 558 758 784 1200 554 524
Usage / MB
I.6 Standard Deviations
We repeat the experiments three times on the HeterNS benchmark and provide standard deviations here. As shown in Table 23-24, Unisolver surpasses the previous state-of-the-art models with high confidence. Note that we compare Unisolver with the second-best model, which is a strong baseline as it is not achieved by a single model. The results demonstrate that Unisolver significantly outperforms baseline models, with the second-best result falling more than three standard deviations behind, except in the case of viscosity $\nu=8e-4$.
@@ I.6 Standard Deviations
I.7 Detailed Compute Resources
@@
Appendix J Full Trajectory Visualizations
To better understand the temporal evolution of the benchmark, we visualize the full trajectory of the ground truth and Unisolver predictions on HeterNS and 2D mixed PDEs in Figure 17 and 18.
Appendix K Limitations and Future Work
This paper presents Unisolver to solve PDEs under universal PDE components, which achieves impressive performance supported by extensive analyses and visualizations. However, our method is currently limited to grid data due to the patchifying process during the embedding of point-wise components. Actually, this limitation is shared in all the generalizable PDE solvers, such as MPP [29], Poseidon [13], PDEformer [45] and DPOT [11]. One fundamental reason is the lack of suitable and large-scale irregular-mesh PDE datasets, which will require extremely high computation costs for generation and massive resources for collection. Since our primary focus in this paper is on the study of model architecture design and generalization capabilities, we would like to leave the irregular-mesh PDE dataset as a future work. Also, the capability to handle irregular meshes of Unisolver can be achieved by replacing the canonical Transformer with the latest geometry-general PDE solver: Transolver [43].